OpenStudy (anonymous):

What is the cofactor of a33=-3 for this matrix: 4 3 1 2 3 -1 1 -1 0 2 -3 0 5 -2 2 1

OpenStudy (nikvist):

cofactor of a33 is determinant of matrix 4 3 2 3 -1 -1 5 -2 1

OpenStudy (anonymous):

So would the answer then be -32?

OpenStudy (nikvist):

4*(-3)-3*8+2*(-1)=-12-24-2=-38